# Realising VCD for untwisted automorphism groups of RAAGs

### v1

@@ 3.3. Bounds on $\textsc{vcd}(U(A_{\Gamma}))$
A thorough analysis of large abelian subgroups of $U(A_{\Gamma})$ is made by Millard and Vogtmann in [MV19], which provides lower bounds on the virtual cohomological dimension of $U(A_{\Gamma})$.
@@
We have $M(L)\leq\textsc{vcd}(U(A_{\Gamma}))\leq dim(K_{\Gamma})=M(V).$
In fact, Millard and Vogtmann go further; they are able to prove that any abelian subgroup generated by $\Gamma$-Whitehead automorphisms has rank at most $M(L)$. Hence, if there is some larger-rank free abelian subgroup of $U(A_{\Gamma})$, then it is not generated by elements of this natural class. We will see later (§6.1) that there are graphs $\Gamma$ for which $M(V)$ is strictly larger than $M(L)$. Thus in these cases, as discussed in Remark 1.1, either these natural large-rank free abelian subgroups generated by $\Gamma$-Whitehead automorphisms fail to realise $\textsc{vcd}(U(A_{\Gamma}))$, or else the dimension of the untwisted spine $K_{\Gamma}$ fails to realise $\textsc{vcd}(U(A_{\Gamma}))$.
@@ 4. The retraction process
Fix a finite simplicial graph $\Gamma$. We now outline a process for performing an equivariant deformation retraction of $K_{\Gamma}$, obtaining a new cube complex which has a proper and cocompact action of $U(A_{\Gamma})$.
@@
We make this intuition of ‘redundancy’ precise in our key lemmas (4.10 & 4.13), which then enable us to perform a $U(A_{\Gamma})$-equivariant retraction of $K_{\Gamma}$ in §4.4. The key lemmas require certain graph theoretic conditions (Definitions 4.8 & 4.11), which means we can only apply our retraction process when $\Gamma$ satisfies these conditions. There is therefore scope to strengthen this technique by improving the notion of ‘redundancy’ to something finer than ‘hugging’. While we have some partial results in this direction, we have decided that ‘hugging’, presented in Definition 4.2, is the best balance between clarity and strength for this document.
@@
For a vertex $v\in V(\Gamma)$, write $\mathcal{C}(v)$ for the set of connected components of $\Gamma\setminus\operatorname{st}({v})$.
@@ 4.1. Hugging partitions
Let $\mathcal{Q}$ be a non-principal partition, based at the vertex $u$. Let $m>_{\circ}u$ be a principal vertex, and let $C\in\mathcal{C}(u)$ be the component which contains $m$. Since $m$ dominates $u$$\mathcal{C}(m)$ is the union of:
@@
    those components of $\Gamma\setminus\operatorname{st}({m})$ which are contained entirely in $C$.
@@
  • One side, say $Q$, contains $C$—and so contains all elements in $\mathcal{C}(m)$ of type (iii), as well as $\{m,m^{-1}\}\cup\left(\operatorname{lk}({m})^{\pm}\setminus\operatorname{lk}({u})^{\pm}\right)$. The rest of the elements of $Q$ (apart from the base element $u$ or $u^{-1}$) are contained in elements of type (i)—that is, the rest of $Q$ is a union of elements of $\mathcal{C}(u)$.
  • The other side, $\overline{Q}$, consists of the remainder of components of type (i) (as well as the base element $u$ or $u^{-1}$).
@@ 4.2. Preparatory lemmas
[…]
If $v_{2}$ and $v_{2}^{-1}$ are in separate components of $\Gamma\setminus\operatorname{st}({v_{1}})$, then  $\operatorname{lk}({v_{1}})\supseteq\operatorname{lk}({v_{2}})$ so $R_{1}\cap\operatorname{lk}({v_{2}})^{\pm}=\emptyset$, and we’re done. So we can suppose that $\{v_{2},v_{2}^{-1}\}$ is contained in one component $C\in\mathcal{C}(v_{1})$.
@@
Remark 4.6.
This lemma appears in [CSV17] as Lemma 3.4; we have rephrased the proof here for ease of comparison.
Note that the statement is symmetric; we can also conclude that $\overline{R_{2}}\subseteq\overline{R_{1}}$.
@@
Claim. $u\neq v$.
Suppose that $\mathcal{R}$ is based at $u$. We have $u\in P_{1}\sqcup P_{2}$; without loss of generality take $u\in P_{1}$. Let $R$ be the side of $\mathcal{R}$ which contains $u$. If $R\subseteq P_{1}$, then $R\subseteq Q$, which contradicts incompatibility of $\mathcal{R}$ and $\mathcal{Q}$. Hence $R$ contains either $P_{1}$ or $\overline{P_{1}}$. We cannot have $R\supseteq\overline{P_{1}}$, as $\overline{P_{1}}\ni u^{-1}$, and $u^{-1}\notin R$ (since $\mathcal{R}$ is based at $u$). Hence $R\supseteq P_{1}$. However, $\mathcal{R}$ cannot split $m$ (since $m$ is principal), so since $m\in R$, we must also have $m^{-1}\in R$. Hence $R$ has non-empty intersection with both sides of $\mathcal{P}_{2}$. Since $u^{-1}\in\overline{R}\cap\overline{P_{2}}$, we must therefore have $\overline{R}\cap P_{2}=\emptyset$ and so $P_{2}\subseteq R$. Hence $R$ contains all of $\mathcal{C}^{Q}(u)$, so $R\supseteq Q$, and once again we have a contradiction to the incompatibility of $\mathcal{R}$ and $\mathcal{Q}$.
This proves the claim, so $u\neq v$. Hence $d_{\Gamma}(u,v)=2$.
@@ 4.3. Key Lemmas
We say that $\Gamma$ satisfies Condition 1 if for every non-principal vertex $u$, and every principal $m>_{\circ}u$, there is no non-principal vertex $u^{\prime}$ with $d_{\Gamma}(u,u^{\prime})=2$ such that $[u^{\prime},m]=1$.
@@
Here, $m$ and $m^{\prime}$ are principal, with $m>_{\circ}u$ and $m^{\prime}>_{\circ}u^{\prime}$ (so $u$ and $u^{\prime}$ are non-principal); we have $d_{\Gamma}(u,u^{\prime})=2$ and $[u^{\prime},m]=1$, so this graph does not satisfy Condition 1. We remark that in this example, the unique vertex $v$ in $\operatorname{lk}({u})$ is principal and commutes with $u^{\prime}$ but is not $>_{\circ}u$, while $a$ is non-principal and satisfies $d_{\Gamma}(u,a)=2$, but does not commute with either of $m$ or $m^{\prime}$ (which are both $>_{\circ}u$).
@@
Denote the base of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ by $m$, and let $C\in\mathcal{C}(u)$ be the component containing $m$.
By the contrapositive of Lemma 4.7, if $u^{\prime}\notin C$, then $\mathcal{Q}^{\prime}$ is compatible with $\mathcal{Q}$. By Lemma 4.4, if $d_{\Gamma}(u,u^{\prime})\notin\{0,2\}$, then $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are compatible. Combining these observations, we may assume that $u^{\prime}\in C$ and that $d_{\Gamma}(u,u^{\prime})=2$. Moreover, since $\Gamma$ satisfies Condition 1, we may assume that $u^{\prime}$ does not commute with $m$ (so in particular, $\mathcal{Q}^{\prime}$ is not adjacent to $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$).
@@
Say that $\Gamma$ satisfies Condition 2 if, for every non-principal vertex $u$, every principal vertex $m>_{\circ}u$, and every principal vertex $n\neq m$ with $d_{\Gamma}(u,n)=2$, we have $[m,n]\neq 1$.
In other words, in every component of $\Gamma\setminus\operatorname{st}({u})$ which contains some principal $m>_{\circ}u$, there is no other principal vertex $n$ commuting with $m$. Intuitively this amounts to connected components of $\Gamma\setminus\operatorname{st}({u})$ not having too many vertices close to $u$, as can be seen in the following example.
@@
Let $\mathcal{R}$ be based at the principal vertex $n$.
Case I: $n\not\sim m$.
If $d_{\Gamma}(u,n)\neq 2$, then by Lemma $\mathcal{R}$ is compatible with $\mathcal{Q}$, so suppose that $d_{\Gamma}(u,n)=2$. Since $\Gamma$ satisfies Condition 2, $[n,m]\neq 1$, so $\mathcal{R}$ and $\mathcal{P}_{1}$ are not adjacent. Nevertheless, $\mathcal{R}$ is compatible with $\mathcal{P}_{1}$, so there must be a choice of sides $P_{1}^{\times},R^{\times}$ such that $P_{1}^{\times}\cap R^{\times}=\emptyset$. If $P_{1}^{\times}=\overline{P_{1}}$, then since $\overline{Q}\subseteq\overline{P_{1}}$ we have $R^{\times}\cap\overline{Q}=\emptyset$, and we’re done. So we may take $P_{1}^{\times}=P_{1}$; let $R$ be the side of $\mathcal{R}$ with $R\cap P_{1}=\emptyset$.
Similarly, we may assume that some side of $\mathcal{R}$ has empty intersection with $P_{2}$. We show that $R\cap P_{2}=\emptyset$. Without loss of generality, assume $R$ is the $n$-side of $\mathcal{R}$. Since $R\cap P_{1}=\emptyset$$n\notin P_{1}$ (and so $n\in\overline{P_{1}}$). By Lemma 4.7, if $n\notin C$$\mathcal{R}$ is automatically compatible with $\mathcal{Q}$, so we may assume that $n\in C$. Hence $n\in Q$, so (since $[n,m]\neq 1$) we have $n\in P_{2}$, so $n\in\overline{P_{1}}\cap P_{2}$. Since $n\not\sim m$ and $n$ is principal, $n$ cannot be split by $\mathcal{P}_{1}$, so $n^{-1}\in\overline{P_{1}}\cap P_{2}$—in particular, we cannot have $\overline{R}\cap P_{2}=\emptyset$.
@@
Note that since $m>_{\circ}u$ and $m>_{\circ}u^{\prime}$, we cannot have $[u,u^{\prime}]=1$: e.g., $u^{\prime}\in\operatorname{lk}({u})\subseteq\operatorname{lk}({m})$, which precludes $m>_{\circ}u^{\prime}$.
@@
Consider the side $Q^{\prime}$ of $\mathcal{Q}^{\prime}$ which contains $m$. Principality of $m$ implies that $\mathcal{Q}^{\prime}$ does not split $m$, so $\{m,m^{-1}\}\subseteq Q^{\prime}$, which means that $Q^{\prime}$ intersects both sides of $\mathcal{P}_{1}$. Hence, since $\mathcal{Q}^{\prime}$ and $\mathcal{P}_{1}$ are compatible, one side of $\mathcal{P}_{1}$ must have empty intersection with $\overline{Q^{\prime}}$. If $\overline{P_{1}}\cap\overline{Q^{\prime}}=\emptyset$, then we have $\overline{Q^{\prime}}\subseteq P_{1}\subseteq R$, and we’re done. So assume that $P_{1}\cap\overline{Q^{\prime}}=\emptyset$. In this case, $Q^{\prime}$ must contain $P_{1}$. Applying the same argument to $P_{2}$, we see that either $\overline{P_{2}}\cap\overline{Q^{\prime}}=\emptyset$ (in which case $\overline{Q^{\prime}}\subseteq P_{2}\subseteq\overline{R}$, and we’re done) or $P_{2}\subseteq Q^{\prime}$—take this to be the case. But now $Q^{\prime}\supseteq\left(P_{1}\cup P_{2}\right)$, which means that
@@
[…]
@@
Let $\Pi$ be a compatible set of partitions, and let $H\subseteq\Pi$ be the set of non-principal partitions in $\Pi$ which are hugged in $\Pi$. Suppose that $\mathcal{R}$ is a principal partition which is compatible with $\Pi\setminus H$. Then $\mathcal{R}$ is incompatible with at most one element of $H$.
Therefore, if $\mathcal{R}$ is incompatible with some $\mathcal{Q}\in H$, then it can replace $\mathcal{Q}$—i.e., $\Pi^{\prime}\coloneqq\left(\Pi\setminus\mathcal{Q}\right)\cup\mathcal{R}$ is a compatible set.
@@
The statement in the lemma follows immediately from the statement here. For the converse, notice that if $\mathcal{R}$ were incompatible with two elements of $H$, then we could take them to be $\mathcal{Q}_{1}$ and $\mathcal{Q}_{2}$ in the lemma, which would be a contradiction. ∎

### v2

@@ 3.3. Bounds on $\textsc{vcd}(U(A_{\Gamma}))$
A thorough analysis of large abelian subgroups of $U(A_{\Gamma})$ is made by Millard and Vogtmann in [MV21], which provides lower bounds on the virtual cohomological dimension of $U(A_{\Gamma})$.
@@
We have $M(L)\leq\textsc{vcd}(U(A_{\Gamma}))\leq\dim(K_{\Gamma})=M(V).$
In fact, Millard and Vogtmann go further; they are able to prove that any abelian subgroup generated by $\Gamma$-Whitehead automorphisms has rank at most $M(L)$. Hence, if there is some larger-rank free abelian subgroup of $U(A_{\Gamma})$, then it is not generated by elements of this natural class. We will see later (§6.1) that there are graphs $\Gamma$ for which $M(V)$ is strictly larger than $M(L)$. Thus in these cases, as discussed after Remark 1.1, either these natural large-rank free abelian subgroups generated by $\Gamma$-Whitehead automorphisms fail to realise $\textsc{vcd}(U(A_{\Gamma}))$, or else the dimension of the untwisted spine $K_{\Gamma}$ fails to realise $\textsc{vcd}(U(A_{\Gamma}))$.
@@ 4. The retraction process
Fix a finite simplicial graph $\Gamma$. We now outline a process for performing a $U(A_{\Gamma})$-equivariant deformation retraction of $K_{\Gamma}$, obtaining a new cube complex which has a proper and cocompact action of $U(A_{\Gamma})$. Since the only complexes we are considering from here on are the untwisted spine $K_{\Gamma}$ and retractions of it, when we say ‘equivariant’ we will always mean with respect to the action of the untwisted subgroup $U(A_{\Gamma})$.
@@
We make this intuition of ‘redundancy’ precise in our key lemmas (4.11 & 4.14), which then enable us to perform a $U(A_{\Gamma})$-equivariant retraction of $K_{\Gamma}$ in §4.4. The key lemmas require certain graph theoretic conditions (Definitions 4.8 & 4.12), which means we can only apply our retraction process when $\Gamma$ satisfies these conditions. There is therefore scope to strengthen this technique by improving the notion of ‘redundancy’ to something finer than ‘hugging’. While we have some partial results in this direction, we have decided that ‘hugging’, presented in Definition 4.2, is the best balance between clarity and strength for this document.
@@
For a vertex $v\in V(\Gamma)$, write $\mathcal{C}(v)$ for the set of connected components of $\Gamma^{\pm}\setminus\operatorname{st}({v})^{\pm}$.
@@ 4.1. Hugging partitions
[…]
@@
    those components of $\Gamma^{\pm}\setminus\operatorname{st}({m})^{\pm}$ which are entirely contained in $C$.
@@
  • One side, say $Q$, contains $C$ – and so contains all elements in $\mathcal{C}(m)$ of type (iii), as well as $\{m,m^{-1}\}\cup\left(\operatorname{lk}({m})^{\pm}\setminus\operatorname{lk}({u})^{\pm}\right)$. The rest of the elements of $Q$ (apart from the base element $u$ or $u^{-1}$) are contained in elements of type (i) – that is, the rest of $Q$ is a union of elements of $\mathcal{C}(u)$.
  • The other side, $\overline{Q}$, consists of the remainder of components of type (i) (as well as the other base element $u$ or $u^{-1}$).
@@ 4.2. Preparatory lemmas
Remark 4.6.
The next lemma appears in [CSV17] as Lemma 3.4; we have rephrased the proof here for ease of comparison.
@@
Since $\mathcal{R}_{1}$ and $\mathcal{R}_{2}$ are compatible and not adjacent, we know that there are sides $R_{1}$$R_{2}$ of $\mathcal{R}_{1}$$\mathcal{R}_{2}$ respectively such that $R_{1}\subseteq R_{2}\cup\operatorname{lk}({v_{2}})^{\pm}$.
If $v_{2}$ and $v_{2}^{-1}$ are in separate components of $\Gamma^{\pm}\setminus\operatorname{st}({v_{1}})^{\pm}$, then  $\operatorname{lk}({v_{1}})\supseteq\operatorname{lk}({v_{2}})$ so $R_{1}\cap\operatorname{lk}({v_{2}})^{\pm}=\emptyset$, and we’re done. So we can suppose that $\{v_{2},v_{2}^{-1}\}$ is contained in one component $C\in\mathcal{C}(v_{1})$.
@@
Note that the statement of the previous lemma is symmetric; we can also conclude $\overline{R_{2}}\subseteq\overline{R_{1}}$.
@@
Claim.
$u\neq v$.
Proof of Claim..
Suppose that $\mathcal{R}$ is based at $u$. We have $u\in P_{1}\sqcup P_{2}$; without loss of generality take $u\in P_{1}$. Let $R$ be the side of $\mathcal{R}$ which contains $u$. If $R\subseteq P_{1}$, then $R\subseteq Q$, which contradicts incompatibility of $\mathcal{R}$ and $\mathcal{Q}$. Hence $R$ contains either $P_{1}$ or $\overline{P_{1}}$. We cannot have $R\supseteq\overline{P_{1}}$, as $\overline{P_{1}}\ni u^{-1}$, and $u^{-1}\notin R$ (since $\mathcal{R}$ is based at $u$). Hence $R\supseteq P_{1}$. However, $\mathcal{R}$ cannot split $m$ (since $m$ is principal), so since $m\in R$, we must also have $m^{-1}\in R$. Hence $R$ has non-empty intersection with both sides of $\mathcal{P}_{2}$. Since $u^{-1}\in\overline{R}\cap\overline{P_{2}}$, we must therefore have $\overline{R}\cap P_{2}=\emptyset$ and so $P_{2}\subseteq R$. Hence $R$ contains all of $\mathcal{C}^{Q}(u)$, so $R\supseteq Q$, and once again we have contradicted the incompatibility of $\mathcal{R}$ and $\mathcal{Q}$. ∎
This proves the claim; $u\neq v$. Hence $d_{\Gamma}(u,v)=2$, as required.
@@ 4.3. Key Lemmas
We say that $\Gamma$ satisfies Condition 1 if for every non-principal vertex $u$, and every principal $m>_{\circ}u$, there is no non-principal vertex $u^{\prime}$ with $d_{\Gamma}(u,u^{\prime})=2$ such that $u^{\prime},m]=1$.
Remark 4.9.
Note that if two vertices vvww satisfy $[v,w]=1$, then we have $d_{\Gamma}(v,w)\in\{0,1\}$ (so in particular, if $v\neq w$, then $v,w]=1\iff d_{\Gamma}(v,w)=1$ (and equivalently, we have $v\in\operatorname{lk}({w})$ and $w\in\operatorname{lk}({v}))$). Hence the end of the preceding definition could be rephrased as “there is no non-principal vertex $u^{\prime}$ with $d_{\Gamma}(u,u^{\prime})=2$ and $d_{\Gamma}(u^{\prime},m)=1$”. However, we will find it useful to interchangeably use both notions throughout the document, so we will use both notations.
@@
Here, $m$ and $m^{\prime}$ are principal, with $m>_{\circ}u$ and $m^{\prime}>_{\circ}u^{\prime}$ (so $u$ and $u^{\prime}$ are non-principal); we have $d_{\Gamma}(u,u^{\prime})=2$ and $u^{\prime},m]=1$, so this graph does not satisfy Condition 1. We remark that in this example, the unique vertex $v$ in $\operatorname{lk}({u})$ is principal and commutes with $u^{\prime}$ but is not $_{\circ}u$, while $a$ is non-principal and satisfies $d_{\Gamma}(u,a)=2$, but does not commute with either of $m$ or $m^{\prime}$ (which are both $_{\circ}u$).
@@
Denote the bases of $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ by $u$ and $u^{\prime}$ respectively, the base of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ by $m$, and let $C\in\mathcal{C}(u)$ be the component containing $m$.
By the contrapositive of Lemma 4.7, if either $u^{\prime}\notin C$ or $d_{\Gamma}(u,u^{\prime})\neq 2$, then $\mathcal{Q}^{\prime}$ is compatible with $\mathcal{Q}$, and we’re done, so assume that $u^{\prime}\in C$ and that $d_{\Gamma}(u,u^{\prime})=2$. Moreover, since $\Gamma$ satisfies Condition 1, we may assume that $u^{\prime}$ does not commute with $m$ (so in particular, $\mathcal{Q}^{\prime}$ is not adjacent to $\mathcal{P}_{1}$ or $\mathcal{P}_{2}$).
@@
Say that $\Gamma$ satisfies Condition 2 if, for every non-principal vertex $u$, every principal vertex $m>_{\circ}u$, and every principal vertex $n\neq m$ with $d_{\Gamma}(u,n)=2$, we have $m,n]\neq 1$.
In other words, in every component of $\Gamma\setminus\operatorname{st}({u})$ which contains some principal $m>_{\circ}u$, there is no other principal vertex $n$ close to $u$ and commuting with $m$. Equivalently, every principal vertex $n$ with $d_{\Gamma}(u,n)=2$ satisfies $d_{\Gamma}(m,n)\geq 2$. Intuitively this amounts to connected components of $\Gamma\setminus\operatorname{st}({u})$ not having too many vertices close to $u$, as can be seen in the following example.
@@
Let $\mathcal{Q}$ be based at $u$, and $\mathcal{Q}^{\prime}$ be based at $u^{\prime}$. Let $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ be based at $m>_{\circ}u$, and let $\mathcal{P}_{1}^{\prime}$ and $\mathcal{P}_{2}^{\prime}$ be based at $m^{\prime}>_{\circ}u^{\prime}$. Let $C\in\mathcal{C}(u)$ be the component containing $m$. Let $Q$ be the side of $\mathcal{Q}$ hugged by the sides $P_{1}$ and $P_{2}$ of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$; take the analogous notation $Q^{\prime}$$P_{1}^{\prime}$$P_{2}^{\prime}$.
Let $\mathcal{R}$ be based at the principal vertex $n$.
If $d_{\Gamma}(u,n)\neq 2$, then by Lemma $4\mathcal{R}$ is compatible with $\mathcal{Q}$, so suppose that $d_{\Gamma}(u,n)=2$.
Case I: $n\not\sim m$.
Since $\Gamma$ satisfies Condition 2, $n,m]\neq 1$, so $\mathcal{R}$ and $\mathcal{P}_{1}$ are not adjacent. Nevertheless, $\mathcal{R}$ is compatible with $\mathcal{P}_{1}$, so there must be a choice of sides $P_{1}^{\times},R^{\times}$ such that $P_{1}^{\times}\cap R^{\times}=\emptyset$. If $P_{1}^{\times}=\overline{P_{1}}$, then since $\overline{Q}\subseteq\overline{P_{1}}$ we have $R^{\times}\cap\overline{Q}=\emptyset$, and we’re done. So we may take $P_{1}^{\times}=P_{1}$; let $R$ be the side of $\mathcal{R}$ with $R\cap P_{1}=\emptyset$.
@@
Note that since $m>_{\circ}u$ and $m>_{\circ}u^{\prime}$, we cannot have $u,u^{\prime}]=1$: otherwise, we would have $u^{\prime}\in\operatorname{lk}({u})\subseteq\operatorname{lk}({m})$, which precludes $m>_{\circ}u^{\prime}$.
@@
Consider the side $Q^{\prime}$ of $\mathcal{Q}^{\prime}$ which contains $m$. Principality of $m$ implies that $\mathcal{Q}^{\prime}$ does not split $m$, so $\{m,m^{-1}\}\subseteq Q^{\prime}$, which means that $Q^{\prime}$ intersects both sides of $\mathcal{P}_{1}$. Hence, since $\mathcal{Q}^{\prime}$ and $\mathcal{P}_{1}$ are compatible, one side of $\mathcal{P}_{1}$ must have empty intersection with $\overline{Q^{\prime}}$. If $\overline{P_{1}}\cap\overline{Q^{\prime}}=\emptyset$, then we have $\overline{Q^{\prime}}\subseteq P_{1}\subseteq R$, and we’re done. So assume that $P_{1}\cap\overline{Q^{\prime}}=\emptyset$. In this case, $Q^{\prime}$ must contain $P_{1}$. Applying the same argument to $P_{2}$, we see that either $\overline{P_{2}}\cap\overline{Q^{\prime}}=\emptyset$ (in which case $\overline{Q^{\prime}}\subseteq P_{2}\subseteq\overline{R}$, and we’re done) or $P_{2}\subseteq Q^{\prime}$ – take this to be the case. But now $Q^{\prime}\supseteq\left(P_{1}\cup P_{2}\right)$, which means that
@@
Suppose, without loss of generality, that $u\in Q$ and $u^{\prime}\in Q^{\prime}$. Then by (2), $u^{\prime}\in Q$, but necessarily $u^{\prime-1}\notin Q$ – as otherwise $u^{\prime-1}\in Q^{\prime}$, which is impossible. Hence $\mathcal{Q}$ splits $u^{\prime}$, which means that $u^{\prime}\leq_{\circ}u$. Switching $u$ and $u^{\prime}$ in this argument gives that $u\leq_{\circ}u^{\prime}$ – hence $\operatorname{lk}({u})=\operatorname{lk}({u^{\prime}})$. But now we have $Q\cap\operatorname{lk}({u^{\prime}})^{\pm}=Q^{\prime}\cap\operatorname{lk}({u})^{\pm}=\emptyset$, so $Q=Q^{\prime}$. This contradicts the hypothesis that $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ are distinct partitions. This finishes the argument. ∎
@@
Suppose that $\Gamma$ satisfies Condition 2. Let $\Pi$ be a compatible set of partitions, and let $H\subseteq\Pi$ be the set of non-principal partitions in $\Pi$ which are hugged in $\Pi$. Suppose that $\mathcal{R}$ is a principal partition which is compatible with $\Pi\setminus H$. Then $\mathcal{R}$ is incompatible with at most one element of $H$.
Therefore, if $\mathcal{R}$ is incompatible with some $\mathcal{Q}\in H$, then it can replace $\mathcal{Q}$ – i.e., $\Pi^{\prime}\coloneqq\left(\Pi\setminus\mathcal{Q}\right)\cup\mathcal{R}$ is a compatible set.
@@
The statement in the lemma follows immediately from the statement here. For the converse, notice that if $\mathcal{R}$ were incompatible with two elements of $H$, then we could take them to be $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ in the lemma, which would be a contradiction. ∎